# ViDoRe V3: A Comprehensive Evaluation of Retrieval Augmented Generation in Complex Real-World Scenarios

### v1

@@ Abstract
António Loisonthanks: Equal contributionthanks: Work done while at Illuin Technology   Quentin Macé1,3   Antoine Edy1 Victor Xing1Tom Balough2Gabriel Moreira2Bo Liu2 Manuel Faysse3Céline Hudelot3Gautier Viaud1 1Illuin Technology  2NVIDIA  3CentraleSupélec, Paris-Saclay {antonio.loison, quentin.mace, antoine.edy}@illuin.tech
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2021) has become the dominant paradigm for knowledge-intensive NLP tasks (Gao et al., 2024; Fan et al., 2024). Yet practical deployments introduce complexities that academic benchmarks often overlook when focusing on single-document textual retrieval. First, documents encode critical information in visual elements such as tables, charts, and images designed for human interpretation, which text-only pipelines often ignore (Abootorabi et al., 2025; Cho et al., 2024b). Second, user queries often require open-ended synthesis, comparison, and reasoning over scattered information, not simple factoid lookup (Tang and Yang, 2024; Thakur et al., 2025; Conti et al., 2025). Third, trustworthy systems must ground responses to specific source locations (e.g., bounding boxes), to mitigate hallucinations (Gao et al., 2023; Ma et al., 2024b).
@@ 2 Related Work
While recent textual RAG benchmarks now capture complex user needs like reasoning or summarizing (Thakur et al., 2025; Tang and Yang, 2024; Su et al., 2024), multimodal evaluation often remains limited to single page queries (Faysse et al., 2025). Multi-page datasets like DUDE (Van Landeghem et al., 2023), M3DocRAG (Cho et al., 2024b), ViDoSeek (Wang et al., 2025) or Real-MM-RAG (Wasserman et al., 2025) prioritize extractive retrieval, lacking the diversity of queries encountered in realistic settings. UniDocBench (Peng et al., 2025) represents a concurrent effort that similarly addresses diverse query types and provides comparative evaluation across multiple RAG paradigms. While this benchmark offers valuable contributions, it relies on synthetically generated queries via knowledge-graph traversal, is restricted to English documents, and constrains grounding annotations to parsed document elements. In contrast, our benchmark offers several complementary strengths: fully human-verified annotations, multilingual coverage, free-form bounding box annotations, and a more systematic evaluation of individual visual RAG pipeline components.
@@ 3.2 Query Generation
We further ensure query diversity by pulling summaries from a heterogeneous set of contexts during the generation process.
Two types of input contexts are used: specific document sections that target local information retrieval and cross-section summaries that target multi-document context retrieval. These summaries are produced through a refined process inspired by ViDoRe V2 macé2025vidorebenchmarkv2raising. First, the text is extracted from PDFs using Docling Auer et al. (2024) along with image descriptions. Then, summaries are generated with Qwen3-235B-Instruct Qwen Team (2025) from each document section. They are clustered to group similar summaries together using Qwen3-Embedding-0.6B Zhang et al. (2025) as embedder, UMAP McInnes et al. (2020) for dimension reduction and HDBSCAN Campello et al. (2013) for clustering. Additionally, cross-section summaries are produced by synthesizing the summaries of 2 to 3 randomly selected sections per cluster. From this pool of summaries, a final subset is curated to maintain a strict balance between single-section and cross-section summaries. The selection also ensures an even distribution across section modalities (text, images, and tables) as defined by the Docling element classification.
@@ 3.3 Answer Detection and Generation
Given the scale of our corpora, manual verification of each page relevance for each query is intractable. We therefore adopt a two-stage annotation pipeline. First, Qwen2.5-VL-32B-Instruct (Bai et al., 2025) pre-filters candidate pages by assessing whether each page image is relevant to the query. Queries whose answers span more than 30 flagged pages are discarded. Human annotators then review the remaining query-page pairs, evaluating query quality and rating page relevance on a three-point scale (Not Relevant, Critically Relevant, Fully Relevant).
@@
For each selected query, annotators were tasked with writing an answer based on the pages they marked as relevant. Given that different annotators might have different answer interpretations and tend not to be exhaustive in their answers, we use Qwen2.5-VL-32B-Instruct to generate a final answer based on the relevant page images marked by the annotators and their answers.
@@
For each relevant page, annotators delineate bounding boxes around content supporting the query and attribute a modality type to each bounding box: Text, Table, Chart, Infographic, Image, Mixed or Other.
Because multiple valid interpretations of bounding boxes can exist, we perform a consistency study to evaluate inter-annotator agreement and establish a human performance upper bound for the task.
@@
Across all 10 datasets, we observe an average IoU of 0.50 and F1 of 0.60. These moderate agreement scores reflect the inherent subjectivity of the task: annotators typically agreed on the relevant content but differed in granularity (Appendix 17), with some marking tight bounds around specific content while others included surrounding context.
@@ 4.1 Retrieval
We analyze ColEmbed-3B-v2, the best-performing retriever we evaluated across query type, content modality, and query language.
@@ 4.2 Final Answer Generation
The correctness of generated answers is assessed against the ground truth final answer by an LLM judge (details in Appendix H). Private datasets are omitted to maintain their integrity.
Some benchmark queries involve general knowledge manageable by LLMs without retrieval. To prevent memorization from confounding our assessment of the RAG pipeline, we stratify queries by difficulty based on parametric knowledge. A query is categorized as easy if any model in a 6-LLM panel answers it correctly without context; otherwise, it is labeled hard. Overall, 48.6 % of queries are easy (see Table˜16 for details).
@@ 4.3 Visual Grounding
Beyond generating correct answers, it is highly desirable for RAG pipelines to identify where in the source documents the answer originates, enabling users to verify the grounding of the query answer. We therefore evaluate the ability of LLMs to generate accurate bounding boxes within their final answer. Among the few LLM families with visual grounding capabilities, we select Qwen3-VL-30B-A3B-Instruct and Gemini 3 Pro for evaluation. For each query, we provide the model with the candidate pages shown to the human annotators and prompt it to answer the query while inserting inline bounding boxes in XML format <bboxes image="N"> ... </bboxes> to delimit relevant content (full instructions in Appendix G).
@@
Inter-annotator agreement on evidence localization reaches an F1 of 0.602, whereas the best-performing models achieve markedly lower scores: 0.089 for Qwen3-VL-30B-A3B-Instruct and 0.065 for Gemini 3 Pro. A page-level analysis (Table˜4) reveals that on pages where humans provided bounding boxes, both models annotated the same page only 16–17 % of the time, while 26–27 % of human-annotated pages received no model annotation at all—highlighting recall as the primary bottleneck. Detailed per-domain results and qualitative analysis appear in Appendix G and 17.
@@ 5 Conclusion
This work introduces ViDoRe V3, a multilingual, human-annotated RAG benchmark that evaluates retrieval, final answer generation, and visual grounding on large industry-relevant document corpora. We design a human-in-the-loop annotation methodology, deployed in a 12,000-hour annotation campaign, that produces diverse realistic queries paired with relevant pages, bounding boxes, and reference answers. Evaluating state-of-the-art RAG pipelines, we find that visual retrievers outperform textual ones, late interaction and textual reranking yield substantial gains, and visual context improves answer generation quality. Looking ahead, ViDoRe V3 highlights several concrete research directions for practical multimodal RAG. Retriever models still struggle on cross-lingual and open-ended queries requiring visual interpretation, while VLMs need improvement in answer generation from multi-page contexts as well as accurate visual grounding. To drive progress in multimodal RAG, ViDoRe V3 has been integrated into the MTEB leaderboard, offering a rigorous framework that fosters the creation of more robust document understanding systems.
@@
Annotator Welfare and Compensation
@@
Data Licensing and Privacy
@@
Linguistic and Geographic Bias
@@ Appendix F ColEmbed-3B-v2 performance breakdown
Appendix G Bounding box annotations
@@
Appendix H Final answer evaluation
@@
Appendix I Visual grounding examples
@@
Appendix J Instructions given to Annotators
@@
Figure˜19 details the step-by-step instructions provided to annotators for assessing page relevance, identifying content modalities, and localizing evidence via bounding boxes. Table˜17 gives the definitions of relevancy scores used by the human annotators.

### v2

@@ Abstract
António Loisonthanks: Equal contributionthanks: Work done while at Illuin Technology   Quentin Macé1,3   Antoine Edy1 Victor Xing1Tom Balough2Gabriel Moreira2Bo Liu2 Manuel Faysse3Céline Hudelot3Gautier Viaud1 1Illuin Technology  2NVIDIA  3CentraleSupélec, Paris-Saclay {quentin.mace, antoine.edy, gautier.viaud}@illuin.techthanks: Contact emails
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2021) has become the dominant paradigm for knowledge-intensive NLP tasks (Gao et al., 2024; Fan et al., 2024). Yet practical deployments introduce complexities that academic benchmarks often overlook when focusing on single-document textual retrieval. First, documents encode critical information in visual elements such as tables, charts, and images designed for human interpretation, which text-only pipelines often ignore (Abootorabi et al., 2025; Cho et al., 2024b). Second, user queries often require open-ended synthesis, comparison, and reasoning over scattered information, not simple factoid lookup (Tang and Yang, 2024; Conti et al., 2025; Thakur et al., 2025). Third, trustworthy systems must ground responses to specific source locations (e.g., bounding boxes), to mitigate hallucinations (Gao et al., 2023; Ma et al., 2024b).
@@ 2 Related Work
While recent textual RAG benchmarks now capture complex user needs like reasoning or summarizing (Thakur et al., 2025; Tang and Yang, 2024; Su et al., 2024), multimodal evaluation often remains limited to single page queries (Faysse et al., 2025). Multi-page datasets like DUDE (Van Landeghem et al., 2023), M3DocRAG (Cho et al., 2024b), ViDoSeek (Wang et al., 2025) or Real-MM-RAG (Wasserman et al., 2025) prioritize extractive retrieval, lacking the diversity of queries encountered in realistic settings. UniDocBench (Peng et al., 2025) represents a concurrent effort that similarly addresses diverse query types and provides comparative evaluation across multiple RAG paradigms. While this benchmark offers valuable contributions, it relies on synthetically generated queries via knowledge-graph traversal, is restricted to English documents, and constrains grounding annotations to parsed document elements. In contrast, our benchmark offers several complementary strengths: fully human-verified annotations, a cross-lingual setup, free-form bounding box annotations, and a more systematic evaluation of individual visual RAG pipeline components.
@@ 3.2 Query Generation
We further ensure query diversity by pulling summaries from a heterogeneous set of contexts during the generation process. Two types of input contexts are used: specific document sections that target local information retrieval and cross-section summaries that target multi-document context retrieval. These summaries are produced through a refined process inspired by ViDoRe V2 macé2025vidorebenchmarkv2raising. First, the text is extracted from PDFs using Docling Auer et al. (2024) along with image descriptions. Then, summaries are generated with Qwen3-235B-Instruct Qwen Team (2025) from each document section. They are clustered to group similar summaries together using Qwen3-Embedding-0.6B Zhang et al. (2025) as embedder, UMAP McInnes et al. (2020) for dimension reduction and HDBSCAN Campello et al. (2013) for clustering. Additionally, cross-section summaries are produced by synthesizing the summaries of 2 to 3 randomly selected sections per cluster. From this pool of summaries, a final subset is curated to maintain a strict balance between single-section and cross-section summaries. The selection also ensures an even distribution across section modalities (text, images, and tables) as defined by the Docling element classification.
@@ 3.3 Answer Detection and Generation
Given the scale of our corpora, manual verification of each page relevance for each query is intractable. We therefore adopt a two-stage annotation pipeline. First, Qwen2.5-VL-32B-Instruct (Bai et al., 2025) pre-filters candidate pages by assessing whether each page image is relevant to the query. Queries whose answers span more than 30 flagged pages are discarded. Human annotators then review the remaining query-page pairs, evaluating query quality and rating page relevance on a three-point scale (Not Relevant, Critically Relevant, Fully Relevant). We selected Qwen2.5-VL-32B-Instruct for its high recall, prioritizing coverage over precision and leaving final relevance judgments entirely to human annotators (see Appendix G for details and distributional validation).
@@
For each selected query, annotators were tasked with writing an answer based on the pages they marked as relevant. Given that different annotators might have different answer interpretations and tend not to be exhaustive in their answers, we use Qwen2.5-VL-32B-Instruct to generate a final answer based on the relevant page images marked by the annotators and their answers. To validate that this aggregation faithfully preserves annotator intent, we evaluated the VLM-aggregated answer against individual annotator responses using a GPT-5.2 judge to assess factual consistency. The aggregated answer matched the exact informational content (with minor paraphrasing) of at least one annotator’s response in 86.3% of cases, confirming the VLM predominantly acts as a selector. For the remaining 13.7% of divergent cases, manual review of a random subset showed the aggregated answer was judged superior in most cases, either by merging complementary information from two incomplete responses or by correcting verifiable factual errors in individual annotations.
@@
For each relevant page, annotators delineate bounding boxes around content supporting the query and attribute a modality type to each bounding box: Text, Table, Chart, Infographic, Image, Mixed or Other. Because multiple valid interpretations of bounding boxes can exist, we perform a consistency study to evaluate inter-annotator agreement and establish a human performance upper bound for the task.
@@
Across all 10 datasets, we observe an average IoU of 0.50 and F1 of 0.60. These moderate agreement scores reflect the inherent subjectivity of the task: annotators typically agreed on the relevant content but differed in granularity (Appendix 17), with some marking tight bounds around specific content while others included surrounding context. Section 4.3 describes how our evaluation methodology accounts for this ambiguity and how model scores should be interpreted relative to this human ceiling.
@@ 4.1 Retrieval
We analyze ColEmbed-3B-v2, the best-performing retriever we evaluated across query type, content modality, and query language. A breakdown by query generation source is provided in Appendix E (Table 14).
@@ 4.2 Final Answer Generation
The correctness of generated answers is assessed against the ground truth final answer by an LLM judge (details in Appendix I). Private datasets are omitted to maintain their integrity.
Some benchmark queries involve general knowledge manageable by LLMs without retrieval. To prevent memorization from confounding our assessment of the RAG pipeline, we stratify queries by difficulty based on parametric knowledge. A query is categorized as easy if any model in a 6-LLM panel answers it correctly without context; otherwise, it is labeled hard. Overall, 48.6 % of queries are easy (see Table˜22 for details).
@@ 4.3 Visual Grounding
Beyond generating correct answers, it is highly desirable for RAG pipelines to identify where in the source documents the answer originates, enabling users to verify the grounding of the query answer. We therefore evaluate the ability of LLMs to generate accurate bounding boxes within their final answer. Among the few LLM families with visual grounding capabilities, we select Qwen3-VL-30B-A3B-Instruct and Gemini 3 Pro for evaluation. For each query, we provide the model with the candidate pages shown to the human annotators and prompt it to answer the query while inserting inline bounding boxes in XML format <bboxes image="N"> ... </bboxes> to delimit relevant content (full instructions in Appendix H).
@@
Inter-annotator agreement on evidence localization reaches an F1 of 0.602, whereas the best-performing models achieve markedly lower scores: 0.089 for Qwen3-VL-30B-A3B-Instruct and 0.065 for Gemini 3 Pro, underlining substantial room for improvement on this task. A page-level analysis (Table˜4) reveals that on pages where humans provided bounding boxes, both models annotated the same page only 16–17 % of the time, while 26–27 % of human-annotated pages received no model annotation at all, highlighting recall as the primary bottleneck. Per-domain results and qualitative analysis appear in Appendix H and 17.
@@ 5 Conclusion
This work introduces ViDoRe V3, a human-annotated RAG benchmark that evaluates cross-lingual retrieval, final answer generation, and visual grounding on large industry-relevant document corpora. We design a human-in-the-loop annotation methodology, deployed in a 12,000-hour annotation campaign, that produces diverse realistic queries paired with relevant pages, bounding boxes, and reference answers. Evaluating state-of-the-art RAG pipelines, we find that visual retrievers outperform textual ones, late interaction and textual reranking yield substantial gains, and visual context improves answer generation quality. Looking ahead, ViDoRe V3 highlights several concrete research directions for practical multimodal RAG. Retriever models still struggle on cross-lingual and open-ended queries requiring visual interpretation, while VLMs need improvement in answer generation from multi-page contexts as well as accurate visual grounding. By providing a rigorous framework for evaluating these limitations, ViDoRe V3 serves as a catalyst for the development of more robust, intelligent document understanding models.
@@
Annotator Welfare and Compensation.
@@
Data Licensing and Privacy.
@@
Linguistic and Geographic Bias.
@@ Appendix F ColEmbed-3B-v2 performance breakdown
Performance by Query Generation Source
Table 14 reports NDCG@10 of ColEmbed-3B-v2 stratified by query type and generation source. Queries generated with full page image access (Human+Image) consistently yield the highest scores, as annotators can anchor queries directly to visible content. Synthetic and human blind-contextual queries (SDG+Summary and Human+Summary) score similarly, with a mean difference of 0.044 NDCG@10, confirming that the synthetic pipeline does not introduce a simplicity bias.
Query Type
SDG+
Sum.
Human+
Sum.
Human+
Img.
Boolean 0.615 0.692 0.801
Compare-Contrast 0.573 0.514 0.743
Open-Ended 0.447 0.459 0.657
Extractive 0.568 0.623 0.743
Enumerative 0.449 0.559 0.697
Numerical 0.592 0.659 0.732
Multi-Hop 0.398 0.378 0.688
Mean 0.521 0.565 0.726
Table 14: ColEmbed-3B-v2 NDCG@10 by query type and generation source. SDG+Sum.: synthetic blind-contextual; Human+Sum:: human blind-contextual; Human+Img:: human with full page image access.
@@
Appendix G VLM Filtering Effect on Query Distribution
Table 16 reports the recall/precision trade-off that motivated the choice of Qwen2.5-VL-32B-Instruct as the pre-filtering model.
Model Recall Precision
Qwen2.5-VL-32B 98% 32%
Gemini-2.5-Flash 78% 81%
Table 16: VLM pre-filtering model selection. Evaluated on 697 human-annotated query–page pairs from ViDoRe v2 (macé2025vidorebenchmarkv2raising) (103 positives, 594 negatives). We selected Qwen2.5-VL-32B as a high-recall sieve to avoid discarding $\sim$25% of relevant documents, leaving final relevance judgments to human annotators.
To assess whether VLM pre-filtering introduces distributional bias, we compare query type, query format, and page modality distributions before and after the filtering step.
Table 17 shows that filtering causes minimal shifts across query types (all $|\Delta|\leq 3\%$), with the exception of Open-Ended queries ($5.4\%$), which reflects the correct removal of overly vague queries. Table 18 shows similarly small shifts across query formats. Table 19 shows that the modality distribution of relevant pages is close to the source documents, confirming that complex visual content is not systematically excluded. The slight decrease in graphical elements use occurs because semantically irrelevant decorative items, such as logos, are naturally not targeted by queries.
Query Type Before (%) After (%) $\Delta$}
Boolean 13.88 16.07 +2.20+2.20
Compare-Contrast 13.73 13.25 0.47-0.47
Enumerative 10.31 10.16 0.15-0.15
Extractive 17.17 20.11 +2.94+2.94
Multi-Hop 8.27 8.06 0.21-0.21
Numerical 6.32 7.42 +1.10+1.10
Open-Ended 30.33 24.93 5.41-5.41
Table 17: Query type distribution before and after VLM filtering.
Query Format Before (%) After (%) $\Delta$}
Instruction 31.32 30.50 0.83-0.83
Keyword 20.00 17.51 2.49-2.49
Question 48.67 51.99 +3.32+3.32
Table 18: Query format distribution before and after VLM filtering.
Source Text Table Graphical
Relevant Pages 55.52% 21.20% 23.28%
Original Docs 54.10% 15.81% 30.09%
Table 19: Modality distribution for relevant pages vs. original documents. The proportion of text, table, and graphical elements is preserved after filtering, confirming that complex visual content is not systematically excluded.
Appendix H Bounding box annotations
@@
Validation on high-agreement pages
To further characterize model grounding performance, we restrict evaluation to query–page pairs with high inter-annotator agreement (mean pairwise IoU $\geq 0.7$). On this subset, Qwen3-VL-30B-A3B-Instruct and Gemini 3 Pro achieve both substantially higher than full-set scores, yet still well below human performance (21). This confirms that visual grounding is a genuine open challenge in current models.
Model Evaluation Set F1 IoU
Qwen3-VL-30B-A3B-Instruct Full Set 0.089
High-Agreement 0.311 0.262
Gemini 3 Pro Full Set 0.065
High-Agreement 0.208 0.155
Table 21: Model grounding performance comparison between the full evaluation set and the high-agreement subset (mean pairwise IoU $\geq 0.7$).
@@
Appendix I Final answer evaluation
@@
Appendix J Visual grounding examples
@@
Appendix K Instructions given to Annotators
@@
Figure˜19 details the step-by-step instructions provided to annotators for assessing page relevance, identifying content modalities, and localizing evidence via bounding boxes. Table˜23 gives the definitions of relevancy scores used by the human annotators.